Solve for y: y2-3y=18

Answer by stanbon(75887)

y2-3y=18
------
y^2 -3y - 18 = 0
------
Factor:
(y-6)(y+3) = 0
---
Ans: y = 6 or y = -3
==========================
Cheers,
Stan H.
===============